OpenStudy (anonymous):

What is the solution set for the open sentence with the given replacement set? 17 – x = 9, {5, 6, 7, 8} A. 5 B. 6 C. 7 D. 8

OpenStudy (anonymous):

in my time they called this "solve for\(x\)"

OpenStudy (welshfella):

yes lol

OpenStudy (anonymous):

\[17-x=9\] solve in two steps a) subtract \(17\) from both sides then b) change the sign of both sides

OpenStudy (anonymous):

this question goes right in the "bad math" pile, to be compiled in a book soon

OpenStudy (anonymous):

?

OpenStudy (anonymous):

17-x=9 17-x=9-17=-8 So, the answer is 8 17-8=9

OpenStudy (anonymous):

thx gigirained

OpenStudy (anonymous):

no probs